Hi! I was wondering if you can help me solve this problem please:
A total of 50 juniors and seniors were given a mathematics test. The 35 juniors attained an average score of 80 while the 15 seniors attained an average of 70 What was the average score for all 50 students who took the test?

Answer by stanbon(48522)

A total of 50 juniors and seniors were given a mathematics test. The 35 juniors attained an average score of 80 while the 15 seniors attained an average of 70 What was the average score for all 50 students who took the test?
-------------
Rule: (sum of scores) divided by (# of students) = average score
So (average*# of students) = sum of scores
-----------------------------------
sum of scores for juniors = 80*35 =2800
sum of scores for seniors = 70*15 =1050
---------
Total sum of scores = 3850
Average for all 50 students = 3850/50 = 77
====================
Cheers,
Stan H.